OpenStudy (anonymous):

consider the follwoing quadratic equation: 56x^2=-19x+15 using the standard form ax^2+bx+c=0 of the quadratic equation above, factor the left hand side of the equation into two linear factors.

myininaya (myininaya):

56x^2+19x-15=0 --------- think: 56(-15)=7(-120)=14(-60)=42(-20)=21(-40)=-21(40) -21+40=19 --------------- so we have 56x^2-21x+40x-15=0 7x(8x-3)+5(8x-3)=0 (7x+5)(8x-3)=0 x=-5/7, x=3/8